The Equation of directrix of a parabola is 3x +4y+15=0 and equation of tangent at vertex is 3x+4y-5=0.Then the length of latus rectum is equal to?

Grade:11

1 Answers

Arun
6 years ago
Distance between tangent at vertex and directrix
 = 15 -(-5)/5
 = 20/5 = 4
 
Which will be equal to a.
 
Hence length of latus rectum = 4a = 4*5 = 20 units
 
 
Regards
Arun (askIITians forum expert)